# A geometric phase approach to quark confinement from stochastic gauge-geometry flows

### v1

@@ Abstract
We apply a stochastic version of the geometric (Ricci) flow, complemented with the stochastic flow of the gauge Yang–Mills sector, in order to seed the chromo-magnetic and chromo-electric vortices that source the area-law for QCD confinement. The area-law is the key signature of quark confinement in Yang–Mills gauge theories with a non-trivial center symmetry. In particular, chromo-magnetic vortices enclosed within the chromo-electric Wilson loops instantiate the area-law asymptotic behaviour of the Wilson loop vacuum expectation values. The stochastic gauge-geometry flow is responsible for the topology changes that induce the appearance of the vortices. When vortices vanish, due to topology changes in the manifolds associated to the hadronic ground states, the evaluation of the Wilson loop yields a dependence on the length of the path, hence reproducing the perimeter law of the hadronic (Higgs) phase of real QCD. Confinement, instead, is naturally achieved within this context as a by-product of the topology change of the manifold over which the dynamics of the Yang–Mills fields is defined. It is then provided by the Aharonov–Bohm effect induced by the concatenation of the compact chromo-electric and chromo-magnetic fluxes originated by the topology changes. The stochastic gauge-geometry flow naturally accomplishes a treatment of the emergence of the vortices and the generation of turbulence effects. Braiding and knotting, resulting from topology changes, namely stochastic fluctuations of the Einstein–Yang–Mills system, stabilise the chromo-magnetic vortices and dynamically induce, as non-trivial topological features, chiral symmetry-breaking. The latter is due to the appearance of torsional degrees of freedom, to which four fermion interactions are related, and to the formation of fermion condensates, in which gluon strings get stabilised by non-trivial braiding.
@@ I Introduction
The plan of the paper is the following. In Sect. II we provide an overview of the center vortex mechanism of confinement, serving as a baseline for further considerations. In Sect. III we outline the (stochastic) gauge-geometry flow formulation and its implication for the dynamics of confinement. In Sect. IV we elaborate on the important details of the out-of-equilibrium dynamics in he thermal time and on the stochastic RF. The formulation of the Wilsonian approach to QCD and dynamical stochastic evolution of topology in the context of confinement is provided in Sect. V. Then, in Sect. VI we present a novel picture of QCD confinement dynamics resorting to the geometric (topology) fluctuations due to the stochastic gauge-geometry flow. Conclusions and outlook follow in Sect. VII. Finally, in Appendix A we provide details of $3$-manifold topology and geometry relevant for our formulation.
@@ III Theoretical framework
IV Out-of-equilibrium dynamics and Renormalization Group
@@
V Wilsonian approach to QCD and topology changes
We can consider the close neighborhood around a Wilson loop, i.e. a solid torus embedded in $M$ with $\gamma$ being the meridian. We can then consider the knot complement $C(\gamma)=S^{3}-\gamma\times D^{2}$. In a dynamical picture, the knot must fluctuate due to the stochastic RF. For small fluctuations, the knot does not change and it is topologically stable. However, for large fluctuations, two parts of the closed curve come close to each other so that the two parts meet and rearrange in a different way. This operation is well-known in knot theory: one gets a new knot which is concordant to the previous knot. There is a result that every knot is changed by the concordance relation to a hyperbolic knot, i.e. a knot where the knot complement admits a hyperbolic geometry.
The scenario in focus is now the following. At low energies one has small fluctuations and the Wilson loop is determined by the closed curve (i.e. a knot) and the loop is determined by the length of the curve. In contrast, for large fluctuations, one has a change of the knot and then one has to consider to corresponding change of the knot induced by the cobordism between the two knots. Then the Wilson loop is determined by the surface (i.e. the cobordism). This behaviour leads to the confinement (note a Wilson’s original argument about the surface behaviour of the Wilson loop). The size of this surface is confined by the facts that every knot is concordant to a hyperbolic knot and the size of surfaces in hyperbolic manifolds (the knot complement) is determined by Mostow rigidity.
It is possible to interpret stochastic term as a stochastic diffeomorphism, i.e. we consider a function $f:M\rightarrow\mathbb{R}$ on the 3-manifold and the RF:
@@ V Wilsonian approach to QCD and topology changes
where $f$ is considered as a generator of diffeomorphisms. This equation can be written as the gradient flow of the functional
@@
Usually, a stochastic term is only a continuous function but every homeomorphism of the 3-manifold can be arbitrarily accurate approximate by a diffeomorphism (topological and smooth 3-manifolds are equal). Alternatively, one can also consider a $C^{2}$ function so that $\partial_{\mu}\partial_{\nu}f$ is only continuous.

### v2

@@ Abstract
We apply a stochastic version of the geometric (Ricci) flow, complemented with the stochastic flow of the gauge Yang–Mills sector, in order to seed the chromo-magnetic and chromo-electric vortices that source the area-law for QCD confinement. The area-law is the key signature of quark confinement in Yang–Mills gauge theories with a non-trivial center symmetry. In particular, chromo-magnetic vortices enclosed within the chromo-electric Wilson loops instantiate the area-law asymptotic behaviour of the Wilson loop vacuum expectation values. The stochastic gauge-geometry flow is responsible for the topology changes that induce the appearance of the vortices. When vortices vanish, due to topology changes in the manifolds associated with the hadronic ground states, the evaluation of the Wilson loop yields a dependence on the length of the path, hence reproducing the perimeter law of the hadronic (Higgs) phase of real QCD. Confinement, instead, is naturally achieved within this context as a byproduct of the topology change of the manifold over which the dynamics of the Yang–Mills fields is defined. It is then provided by the Aharonov–Bohm effect induced by the concatenation of the compact chromo-electric and chromo-magnetic fluxes originated by the topology changes. The stochastic gauge-geometry flow naturally accomplishes a treatment of the emergence of the vortices and the generation of turbulence effects. Braiding and knotting, resulting from topology changes, namely stochastic fluctuations of the Einstein–Yang–Mills system, stabilise the chromo-magnetic vortices and dynamically induce, as non-trivial topological features, chiral symmetry-breaking. Finally, we observe that dimensional transmutation for the Yang-Mills fields can be derived from the scaling property of the geometric part of the stochastic flow. Specifically, a relation can be derived that involves the infrared equilibrium limit of the Planck constant and yields the correct order of magnitude for $\Lambda_{\rm QCD}$.
@@ I Introduction
The plan of the paper is the following. In Sect. II we provide an overview of the center vortex mechanism of confinement, serving as a baseline for further considerations. In Sect. III we outline the (stochastic) gauge-geometry flow formulation and its implication for the dynamics of confinement. In Sect. V we elaborate on the important details of the out-of-equilibrium dynamics in he thermal time and on the stochastic RF. The formulation of the Wilsonian approach to QCD and dynamical stochastic evolution of topology in the context of confinement is provided in Sect. VI. Then, in Sect. VIII we present a novel picture of QCD confinement dynamics resorting to the geometric (topology) fluctuations due to the stochastic gauge-geometry flow. Conclusions and outlook follow in Sect. IX. Finally, in Appendix A we provide details of $3$-manifold topology and geometry relevant for our formulation.
@@ III Theoretical framework
An example of changing topology induced by the stochastic geometric flow arises as follows. It is known that the Ricci flow can produce singularities when the scalar curvature is negative Perelman (2002); Kleiner and Lott (2008), and this is dealt with by means of surgery Perelman (2003). The idea of the Ricci flow with surgery is to remove the singularity fibers of a product of type $(-c,c)\times\mathbb{S}^{3}$ where the singularity is forming, and cap off $\{-c,c\}\times\mathbb{S}^{3}$. We therefore have (at finite time $T$) the formation of a neck-pinch singularity and we can find a sequence of times $t_{n}$, where $t_{n}\longrightarrow T^{-}$ such that $\lim_{n}|{\rm Rm}|\longrightarrow+\infty$, where $|{\rm Rm}|$ is the largest absolute value of the sectional curvature. In the stochastic geometric flow, the expectation values of the corresponding absolute values of sectional curvatures diverge as well, and singularities can form. In fact, a standard application of Ito’s Lemma gives that the expectation value of $X_{t}\tau$ during an Ito process, as in the stochastic geometric flow (Lulli et al. (2022)), evolves over time according to the equality
$\lim_{\Delta\tau\rightarrow 0}\frac{E_{\tau+\Delta\tau}(X_{\tau+\Delta\tau})-E_{\tau}(X_{\tau})}{\Delta\tau}=\mu_{t}(X_{t}),$ (42)
where $dX_{\tau}=\mu_{\tau}(X_{\tau})d\tau+\sigma_{\tau}(X_{\tau})dW_{\tau}$. In the stochastic geometric flow it means that the expectation values for the metric $g$ evolve as in the Ricci flow. Therefore, the unboundedness of $|{\rm Rm}|$ implies that the expectation value of $|{\rm Rm}|$ is unbounded as well during the stochastic geometric flow, which corresponds to the development of a singularity in finite time. We assume that at the formation of singularities a surgery cobordism removes the neck-pinch. However, we do not restrict ourselves to capping off with balls, which implies that the topology of the manifold changes. The simplest surgery is the attachment of two $1$-handles $D^{1}\times D^{3}$ to two pairs copies of $\mathbb{S}^{3}$, corresponding to collapsed fibers for the formation of two different singularities. This process changes the fundamental group of the base manifold, introducing nontrivial loops wrapping around the attached handles. This nontrivial elements of the fundamental group $\pi_{1}(M)$, where $M$ is the ground manifold, correspond via the standard surjection $\pi_{1}(M)\longrightarrow\frac{{\rm Hol}_{A}}{{\rm Hol^{0}}_{A}}$ for a connection $A$. These nontrivial homotopy classes of $\pi_{1}(M)$ correspond to Wilson loops that wrap around vortexes, whose creation is due to topological fluctuations induced by the creation of singularities in the stochastic geometric flow.
We will later analyse this topology changing process. Instead of attaching two 1-handles we will use a pair of one 1-handle and one 2-handle. The details can be found below.
@@
IV Topological defects induced by the gauge flow
The realisation of this program enforces us to follow a slightly different approach. We now investigate the spatial component of the spacetime, the 3-manifold, and its variation of topology. At this purpose, we modify the Langevin equations (35) and (36) to be
$\!\!\!\!\frac{\partial A^{a}_{m}}{{\partial}\tau}(x^{\prime}\!,\!\tau)\!=$ $\!\!\!\!\!\!\!\epsilon_{m}^{nl}F^{a}_{nl}(x^{\prime}\!,\!\tau)+\xi^{a}_{m}(x^{\prime}\!,\!\tau)\,,$ (43)
$\!\!\!\!\frac{\partial g_{mn}}{{\partial}\tau}(x^{\prime}\!,\!\tau)\!=$ $\!\!\!\!\!\!\!\frac{1}{2\kappa}\!\Big{[}R_{mn}(x^{\prime}\!,\!\tau)\!-\!\nabla_{m}\nabla_{n}f\Big{]}\!\!+\!\xi_{mn}(x^{\prime}\!,\!\tau)\,,$ (44)
where $x^{\prime}=(x^{1},x^{2},x^{3})$ denotes the spatial coordinates, $\tau$ is the stochastic time, and for the indices $l,m,n=1,2,3$. The epsilon tensor is only nonzero for $\epsilon_{1}^{23}=1$ and cyclic permutation, being antisymmetric in the exchange of indices. The function $f$ is the generator of the diffeomorphism for the 3-manifold $\Sigma$. Both equations are the gradient flow for a corresponding functional. In Eq. (43) equation, $f$ introduces the Chern-Simons functional, while in Eq. (44) $f$ introduces the entropy functional of Perelman. We will discuss in what follows a 4D version of this functional. According to Ref. Floer (1988), Eq. (43) is equivalent to the instanton equation. This can be shown by identifying the 0th component of the strength tensor with $F_{0\mu}=dA_{\mu}/dt$, where the derivative with respect to the time coordinate $t$ is expressed as a covariant derivative. We then obtain
$F_{\mu\nu}=\epsilon_{\mu\nu}^{\rho\kappa}F_{\rho\kappa}+\xi_{\mu\nu}\,,$ (45)
which is the stochastic version of the instanton equation, with the epsilon tensor extended to 4D, introducing $\epsilon_{01}$. Eq. (45) is important for several reasons. As shown by Seiberg Seiberg (1988), the instanton solutions dominate the non-perturbative regime of a supersymmetric version of QCD. Furthermore, the theory is induced by a gradient flow of a Chern-Simons functional on the 3-manifold, which is known to induce vortices and its condensation. More importantly, the stochastic term can change the topological class of the theory, as we proceed to show now.
In the following discussion, we switch for compactness of the notation to the language of differential forms. Then the stochastic instanton equation is given by
$F=\star F+\xi\,.$ (46)
The minimum of the Yang-Mills functional
$\int F\wedge\star F$
is the topological invariant (second Chern class),
$c_{2}=\frac{1}{8\pi^{2}}\int F\wedge F\,,$
i.e. the solution of the instanton previously reported.
We may now discuss the singularities of the gauge flow. As an example, we start considering the case of electromagnetism, i.e. a $U(1)$ gauge theory. The gauge potential $A$ with field strength $F=dA$ has to fulfill the equations
$dF=0\,,\qquad d*F=*j\,,$
the 1-form $j$ denoting the source. The equation $dF=d^{2}A=0$ is a direct consequence of the relation $d^{2}=0$. In presence of a monopole, the equation must be modified into
$dF=j_{m}\,,$ (47)
$j_{m}$ denoting the current of the magnetic charges. Eq. (47) can be fulfilled by singular currents, according to the relation (in complex analysis)
$d\left(\frac{dz}{z}\right)=d^{2}(ln(z))=2\pi\imath\delta(z)dz\wedge d\bar{z}\,.$
Nonetheless, this relation can be also fulfilled by a non-trivial complex line bundle (or a non-trivial $U(1)$ bundle) over the 2-sphere $S^{2}$ (or equivalently on $\mathbb{R}^{3}\setminus 0$). Thus, the singularity can be understood as originating from a map between a trivial and a non-trivial complex line bundle. If we consider the stochastic gauge flow for a $U(1)$ gauge theory, then a singularity in this flow will produce a non-trivial bundle, or a monopole.
We may now extend this idea to the stochastic gauge flow of a $SU(3)$ gauge theory. Singularities of this flow are singular bundle maps producing non-trivial $SU(3)$ bundles that are classified by the instanton number. We start focusing on the definition of a differential $df$ of a singular map $f$, using the work by Harvey and Lawson in Ref. Harvey and Lawson (1993), and first describe the general situation.
Let $F$ and $E$ be two vector bundles over $M$. Furthermore, let $\alpha:E\rightarrow F$ be a bundle map admitting singularities, i.e. a subset $\Sigma\subset M$ where the (local) map $\alpha_{x}:E_{x}\rightarrow F_{x}$ is not injective. Let $D_{E}$ and $D_{F}$ be the connections on $E$ and $F$, respectively. The main problem is now that the inverse of the map $\alpha$ not always exists. Instead, we have to define a map $\beta:F\rightarrow E$. At first we assume such a map to exist. Then we define the push forward connection $\mathbf{D}^{\alpha}$ by
$\mathbf{D}^{\alpha}=\alpha\circ D_{E}\circ\beta+D_{F}\circ(1-\alpha\beta)\,.$
On the two bundles, we may introduce metrics that allow to define the adjoint $\alpha^{*}$ of $\alpha$ via the scalar product. Then, outside the singular set $\Sigma$, we define
$\beta=(\alpha^{*}\alpha)^{-1}\alpha^{*}\,.$
In general this procedure breaks down on the singular set $\Sigma$, since $\beta$ becomes singular on $\Sigma$. To amend this situation, we choose an approximation mode, i.e. a fixed smooth function $\chi:[0,\infty]\rightarrow[0,1]$, with $\chi^{\prime}\leq 0,\chi(0)=0$ and $\chi(\infty)=1$. We then define a smooth approximation $\beta_{s}$ to $\beta$ by mean of the relation
$\beta_{s}=\chi\left(\frac{\alpha^{*}\alpha}{s}\right)\beta\,,$
with $s$ some real positive parameter. The family of maps $\beta_{s}$ defines a family of connections $\mathbf{D}_{s}^{\alpha}$ on $F$. As $s\rightarrow 0$, the map $\beta_{s}$ converges to $\beta$ uniformly on a compact subset in $M-\Sigma$. The family $\beta_{s}$ of maps converges for $s\rightarrow\infty$ pointwise to a connection on $F$ for all points in $M$ — see e.g. Ref.  Harvey and Lawson (1993) for proofs.
Within this general context, we can define the connection change by setting $E$ and $F$ to the trivial and non-trivial $SU(3)$ bundle, and $\alpha=df$ to represent the defect induced by the gauge flow. Then the substitute for the inverse $df^{-1}$ is defined by the limit $s\rightarrow\infty$ of $\beta_{s}$ constructed above.
We may now apply this theory to the singularities of the gauge flow. Let $E_{0}$ be a trivial $SU(3)$ bundle over the 4-sphere, regarded as a compactified $\mathbb{R}^{4}$. We can also consider the Minkowski space $\mathbb{R}^{3,1}$, with compactification $S^{3}\times S^{1}$. We may now consider the bundle $E_{0}$ over the spacetime $M$. According to the theory of Harvey and Lawson, every singularity of the gauge flow is related to a (singular) bundle map $a:E_{0}\rightarrow E_{1}$ from the trivial bundle $E_{0}$ to the non-trivial bundle $E_{1}$. The corresponding invariant is the second Chern class
$c_{2}(E_{1})=\frac{1}{8\pi^{2}}\intop_{M}tr(F\wedge F)\,,\qquad c_{2}(E_{0})=0$
for the curvature $F$ of the bundle $E_{1}$.
According to this procedure, we may derive the curvature
$F=d\left(g_{a}^{-1}dg_{a}\right)$ (48)
induced by the singular gauge transformation $g_{a}$ from the bundle map $a:E_{0}\rightarrow E_{1}$. The expression for the curvature represents the singular part of $d\left(g_{a}^{-1}dg_{a}\right)$, the other part $g_{a}^{-1}dg_{a}\wedge g_{a}^{-1}dg_{a}$ being neglected according to the Maurer-Cartan equation.
This formula has two tremendous consequences. Every singularity produces a non-trivial instanton — i.e. a solution of the equation $F=\pm*F$. Then there is a simple relation between the holonomy along a closed loop $\gamma$ around the singularity generated by
[…]
with $A=g_{a}^{-1}dg_{a}$. The integral over the curvature $F$, calculated according to (48), is generated by singularity, having integrated over a surface $S$, with boundary $\gamma$. The surface $S$ always exists, and is known as the Seifert surface. The loop $\gamma$ cannot be contracted, thus the integral over the surface does not vanish. Unfortunately, this is still not enough to derive the scaling behaviour of the integral. We also require the stochastic geometry (Ricci) flow.
We then produce an instanton for every singularity. Every solution $F$ also induces a unique solution $*F$. Consequently, for every color-electric loop (or Wilson loop), there must exist a color-magnetic loop (or t’Hooft loop), where both loops are linked.
V Out-of-equilibrium dynamics and Renormalization Group
@@
VI Wilsonian approach to QCD and topology changes by the stochastic geometry flow
In this section we will analyse the stochastic geometry (Ricci) flow and its implications for the topology of the Wilson loop. In section IV we discussed the appearance of Wilson loops and t’Hooft loops due to the singularities of the gauge flow. The gauge flow is coupled to the geometry flow. Then, the stochastic geometry flow induces fluctuations of the Wilson loops by the fluctuating metric of the embedding space of the loop. We may now discuss the possible fluctuations of the Wilson loops. At first, we clarify the notation. The Wilson loop is an unknotted solid torus: we use the word loop or unknot to express the unknotness. The name “Wilson knot” will be used in the following to indicate a knotted Wilson loop represented by a knotted solid torus. In a dynamical picture, the Wilson knot must fluctuate due to the stochastic geometry flow. For small fluctuations, the Wilson knot does not change and it is topologically stable. However, for large fluctuations, two parts of the closed curve come close to each other, so that the two parts meet and rearrange in a different way. Then a trivial Wilson knot (i.e. a Wilson loop) changed to a knotted Wilson loop (i.e. a Wilson knot). This operation is well-known in knot theory: one obtains a new knot that is concordant to the previous knot. There is a result that every knot is changed by the concordance relation to a hyperbolic knot, i.e. a knot where the knot complement defined below admits a hyperbolic geometry Myers (1983).
We may now consider the close neighborhood around a Wilson loop, i.e. a solid torus embedded in $M$, with $\gamma$ being its meridian. We can then consider the knot complement $C(\gamma)=S^{3}-\gamma\times D^{2}$. This represents the geometry around the knot, i.e. how the knot has an impact on the geometry of the space.
This brings to the following scenario. At low energies one has small fluctuations. The Wilson loop is determined by the unknotted, closed curve (i.e. the unknot or loop), and the loop is supported by the length of the curve. In contrast, large fluctuations induce a change of the Wilson knot, so that it is necessary to consider the corresponding change of the knot induced by the cobordism between the two knots. Then the Wilson knot is determined by the surface, i.e. the cobordism. This behavior leads to the confinement Wilson (1974). The size of this surface is constrained by the fact that every knot is concordant to a hyperbolic knot, thus the size of surfaces in hyperbolic manifolds (the knot complement) is constrained by Mostow rigidity. This is the outline of what we specify below.
We first provide a possible interpretation of the stochastic noise as a stochastic diffeomorphism. A this purpose, we consider a function $f:M\rightarrow\mathbb{R}$ on the 3-manifold and the geometry flow defined by
@@ VI Wilsonian approach to QCD and topology changes by the stochastic geometry flow
where $f$ is considered as the generator of diffeomorphisms. This equation can be written as the gradient flow of the functional
@@
where $f$ can then be interpreted as dilaton. Usually, a stochastic term is only a continuous function. Nonetheless, every homeomorphism of the 3-manifold can be approximated with arbitrary accuracy by a diffeomorphism (topological and smooth 3-manifolds are equal). Alternatively, one can also consider a $C^{2}$ function, in such a manner that $\partial_{\mu}\partial_{\nu}f$ is only continuous.